OpenStudy (anonymous):

Evaluate the integral e^tanx sec^2 (x) dx

OpenStudy (anonymous):

u = tanx can you go from here?

OpenStudy (anonymous):

What is the approximate circumference of a circle with diameter fourteen centimeters? Use twenty-two sevenths to approximate pi. eleven centimeters twenty-two centimeters thirty-three centimeters forty-four centimeters please help me

OpenStudy (johnweldon1993):

As sourwing has stated...if you do a simple u-sub making u = tan(x)...the derivative of tan(x) = ?

OpenStudy (anonymous):

Yeah I got the far, my final answer is e^tanx+c but I'm not sure if it's right

OpenStudy (johnweldon1993):

That is correct! @TheAni-Trix

OpenStudy (johnweldon1993):

\[\large \int\limits^{}_{}e^udu = e^u + c = e^{\tan(x)} + c\]

OpenStudy (anonymous):

Oh okay thank you. ^- ^

OpenStudy (johnweldon1993):

No problem :)